# Quantum Steenrod operations of symplectic resolutions

### v2

@@ 5.1. Equivalence of $p$-curvature and quantum Steenrod operations
Take $X=T^{*}(G/B)$ with its action of ${\mathbf{T}}=T\times S^{1}$ from the maximal torus $T$ and the rotation $S^{1}$ of cotangent fibers. Fix $b\in H^{2}(X;\mathbb{F}_{p})$. We consider the ${\mathbf{T}}$-equivariant quantum connection $\nabla_{b}$ and the ${\mathbf{T}}$-equivariant quantum Steenrod operations $\Sigma_{b}$. Denote by $F_{b}=\nabla_{b}^{p}-t^{p-1}\nabla_{b}$ the $p$-curvature in the direction of $b$.
Recall that $t,\lambda,h$ represent the equivariant parameters for $\mathbb{Z}/p$-action on the source curve $C$ (acting trivially on $X$), the action of torus $T\leq{\mathbf{T}}$ on $X$ (preserving the holomorphic symplectic form), and the action of $S^{1}\leq{\mathbf{T}}$ on $X$ (scaling the holomorphic symplectic form), respectively.
Given the properties of the quantum Steenrod operations $\Sigma_{b}$ established previously, the following proof follows the structure of the proof of 4.8 due to [EV24a].
@@
Consider both operators $F_{b}$ and $\Sigma_{b}$ as endomorphisms of $H^{*}_{\mathbf{T}}(X;\Lambda)_{\mathrm{loc}}{[\![t,\theta]\!]}$ of degree $2p$. Fix the basis (as a $H^{*}_{{\mathbb{Z}/p}\times{\mathbf{T}}}(\mathrm{pt};\Lambda)_{\mathrm{loc}}$-module) given by the stable envelopes, see 3.7. Since the torus fixed locus $X^{T}$ are isolated points, the stable basis is homogeneous in cohomological degrees; $\mathrm{Stab}_{+}(w)\in H^{\dim_{\mathbb{C}}X}_{{\mathbf{T}}}(X)$ for all $w\in W$. Hence the structure constants of both operators are given by expressions of cohomological degree $2p$, which are polynomials in the equivariant parameters $t,\lambda,h$ and power series in the Novikov parameters $q^{A}$.
@@
(5.3) $b\ \ast_{\ {\mathbf{T}}}=b\smile_{{\mathbf{T}}}+b\ \ast_{\mathrm{quantum},{\mathbf{T}}}=(hB^{(0)}_{\mathrm{cl}}+B^{(1)}_{\mathrm{cl}})+hB_{q}^{(0)}$
where $B^{(0)}_{\mathrm{cl}}$, $B^{(1)}_{\mathrm{cl}}$, $B^{(0)}_{q}$ have entries of cohomological degrees $0,2,0$ respectively. Note that $B^{(1)}_{\mathrm{cl}}$ is simply the matrix of $b\smile_{T}$ in the equivariant cohomology for the smaller torus $T\leq{\mathbf{T}}$.
First note that $F_{b}^{(p)}=\Sigma_{b}^{(p)}$ from 2.18 and 4.5, as both matrices are equal to the matrix of the operator $\mathrm{St}^{T}(b)\smile_{T}$.
Also note that $F_{b}^{(0)}=\Sigma_{b}^{(0)}$, as both matrices are equal to the matrix of the $h^{p}$-coefficient of the operator $\overbrace{b\ast_{\mathbf{T}}\cdots\ast_{\mathbf{T}}b}^{p}\ast_{\mathbf{T}}$ for the following degree reasons. For $\Sigma_{b}$, note that $\Sigma_{b}=\overbrace{b\ast_{\mathbf{T}}\cdots\ast_{\mathbf{T}}b}^{p}\ast_{\mathbf{T}}+O(t)$ so the only terms divisible by $h^{p}$ must occur from the $p$-fold quantum multiplication. For $F_{b}=\nabla_{b}^{p}-t^{p-1}\nabla_{b}$, the same is true as can be seen from the expression $\nabla_{b}=t\partial_{b}+b\ast_{{\mathbf{T}}}$.
@@
(5.4) $F_{b}-\Sigma_{b}=h^{p-1}\left(F_{b}^{(1)}-\Sigma_{b}^{(1)}\right)+\cdots+h\left(F_{b}^{(p-1)}-\Sigma_{b}^{(p-1)}\right),$
and the characteristic polynomial of this matrix
(5.5) $\chi(F_{b}-\Sigma_{b})=\sum_{i=0}^{n}\mathrm{tr}\left(\Lambda^{i}(F_{b}-\Sigma_{b})\right)x^{n-i}$
satisfies the following properties: for each $1\leq i\leq n$,
@@
Since $F_{b}$ and $\Sigma_{b}$ both commute with the shift operators from 4.5 and 2.31, so is the difference $F_{b}-\Sigma_{b}$. Consider the shift operator $S_{\beta}$ for the cocharacter $\beta:S^{1}\to{\mathbf{T}}$ that vanishes on all $T$-weights and is weight $1$ for the $S^{1}$-action rotating the holomorphic symplectic form. By the invertibility of shift operators 3.21, we can conjugate $F_{b}-\Sigma_{b}$ by the shift operator which shows the invariance of the characteristic polynomial $\chi(F_{b}-\Sigma_{b})$ under shift of equivariant parameters $h\mapsto h-t$.
@@
For almost all $p$, the $p$-curvature and the quantum Steenrod operation for $b\in H^{2}(X;\mathbb{F}_{p})$ agree.
@@
From 5.1, we know that the difference $F_{b}-\Sigma_{b}$ is a nilpotent operator acting on $H^{*}_{{\mathbf{T}}}(X;\Lambda)_{\mathrm{loc}}{[\![t,\theta]\!]}$ as a $H^{*}_{{\mathbb{Z}/p}\times{\mathbf{T}}}(\mathrm{pt})_{\mathrm{loc}}$-module. From 2.20, $\Sigma_{b}$ commutes with $\nabla_{b}$ and hence with $F_{b}=\nabla_{b}^{p}-t^{p-1}\nabla_{b}$. In particular, it suffices that $F_{b}$ has simple spectrum to show that $F_{b}=\Sigma_{b}$.
To see that the eigenvalues of $F_{b}$ are all distinct, consider the discriminant $D(F_{b})$. Since the equivariant parameter $t\in H^{2}_{\mathbb{Z}/p}(\mathrm{pt})$ is not inverted, one can consider the specialization $D(F_{b})|_{t=0}=D(b^{\ast_{\mathbf{T}}p}\ast_{\mathbf{T}})$; if the operator $b^{\ast_{\mathbf{T}}p}\ast_{\mathbf{T}}$ has simple spectrum, then the discriminant of it would be nonzero, and $D(F_{b})$ would also be nonzero. Therefore it suffices to check that $b^{\ast_{\mathbf{T}}p}\ast_{\mathbf{T}}$ has simple spectrum; this follows if $b{\ast_{\mathbf{T}}}$ has simple spectrum.
Consider the eigenvalues of $b\ast_{\mathbf{T}}|_{q^{A}=0}=b\smile_{\mathbf{T}}$, which are the weights of the line bundle corresponding to $b$ restricted to the $T$-fixed points of $X$. These weights are indeed distinct in $H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p})$ for $X=T^{*}(G/B)$, for all but finitely many primes $p$. Hence $b\smile_{{\mathbf{T}}}$ has simple spectrum, which implies that $b\ast_{\mathbf{T}}$ has simple spectrum. ∎
@@
Proceeding as in Section 3.4, this result also applies to further examples including $X=T^{*}(G/P)$ (the generalized Springer resolutions) or quiver varieties of type $A$, which have (i) isolated $X^{T}$ and hence a stable basis of homogeneous cohomological degree, (ii) simple spectrum of the classical (hence also quantum) multiplication operator. The Springer resolution $T^{*}(G/B)$ in type $A$ are in the intersection of these two classes of examples.

### v3

@@ 5.2. Equivalence of $p$-curvature and quantum Steenrod operations
Take $X=T^{*}(G/B)$ with its action of $\mathbf{T}=T\times S^{1}$ from the maximal torus $T$ and the rotation $S^{1}$ of cotangent fibers. Fix $b\in H^{2}_{\mathbf{T}}(X;\mathbb{F}_{p})$. We consider the $\mathbf{T}$-equivariant quantum connection $\nabla_{b}$ and the $\mathbf{T}$-equivariant quantum Steenrod operations $\Sigma_{b}^{\mathbf{T}}$. Denote by $F_{b}=\nabla_{b}^{p}-t^{p-1}\nabla_{b}$ the $p$-curvature in the direction of $b$.
For this section, with the $\mathbf{T}$-equivariance understood, we will simply denote $\Sigma_{b}^{\mathbf{T}}$ by $\Sigma_{b}$.
Recall that $t,\lambda,h$ represent the equivariant parameters for $\mathbb{Z}/p$-action on the source curve $C$ (acting trivially on $X$), the action of torus $T\leq\mathbf{T}$ on $X$ (preserving the holomorphic symplectic form), and the action of $S^{1}\leq\mathbf{T}$ on $X$ (rotating the holomorphic symplectic form), respectively.
Given the properties of the quantum Steenrod operations $\Sigma_{b}$ established previously, the following proof follows the structure of the proof of 4.9 due to [EV24a].
@@
Consider both operators $F_{b}$ and $\Sigma_{b}$ as endomorphisms of $H^{*}_{\mathbf{T}}(X;\Lambda)_{\mathrm{loc}}{[\![t,\theta]\!]}$ of degree $2p$. Fix the basis of $H^{*}_{\mathbf{T}}(X;\Lambda)_{\mathrm{loc}}{[\![t,\theta]\!]}$ (as a $H^{*}_{{\mathbb{Z}/p}\times\mathbf{T}}(\mathrm{pt};\Lambda)_{\mathrm{loc}}=\Lambda(\!(\lambda,h)\!)[\![t,\theta]\!]$-module) given by the stable envelopes, see 3.8. Since the torus fixed locus $X^{T}$ are isolated points, the stable basis is homogeneous in cohomological degrees; that is, $\mathrm{Stab}_{+}(w)\in H^{\dim_{\mathbb{C}}X}_{\mathbf{T}}(X)$ for all $w\in W$.
Therefore the structure constants of both operators in the stable basis are given by expressions of cohomological degree $2p$, which are polynomials in the equivariant parameters $t,\lambda,h$ and power series in the Novikov parameters $q^{A}$. Polynomiality follows from that $|q^{A}|=0$ from $c_{1}=0$, and moreover that the regularity of the relevant moduli spaces ensure that the equivariant parameters only appear with non-negative powers.
@@
(5.3) $b\ \ast_{\ \mathbf{T}}=b\cup_{\mathbf{T}}+b\ \ast_{\mathrm{quantum},\mathbf{T}}=(hB^{(0)}_{\mathrm{cl}}+B^{(1)}_{\mathrm{cl}})+hB_{q}^{(0)}$
where $B^{(0)}_{\mathrm{cl}}$, $B^{(1)}_{\mathrm{cl}}$, $B^{(0)}_{q}$ have entries of cohomological degrees $0,2,0$ respectively. Note that $B^{(1)}_{\mathrm{cl}}$ is simply the matrix of $b\cup_{T}$ in the equivariant cohomology for the smaller torus $T\leq\mathbf{T}$.
First note that $F_{b}^{(p)}=\Sigma_{b}^{(p)}$ from 2.18 and 4.6, as both matrices are equal to the matrix of the operator $\mathrm{St}^{T}(b)\cup_{T}$. This is due to the fact that $h=0$ limit corresponds to considering counts non-equivariantly with respect to the dilating $S^{1}$-action, for which the Gromov–Witten theory is trivial (cf. [BMO11, Section 1.3]).
Also note that $F_{b}^{(0)}=\Sigma_{b}^{(0)}$, as both matrices are equal to the matrix of the $h^{p}$-coefficient of the operator $\overbrace{b\ast_{\mathbf{T}}\cdots\ast_{\mathbf{T}}b}^{p}\ast_{\mathbf{T}}$ for the following degree reasons. For $\Sigma_{b}$, note that $\Sigma_{b}=\overbrace{b\ast_{\mathbf{T}}\cdots\ast_{\mathbf{T}}b}^{p}\ast_{\mathbf{T}}+O(t)$ so the only terms divisible by $h^{p}$ must occur from the $p$-fold quantum multiplication. For $F_{b}=\nabla_{b}^{p}-t^{p-1}\nabla_{b}$, that the same is true can be seen from the expression $\nabla_{b}=t\partial_{b}+b\ast_{\mathbf{T}}$.
@@
(5.4) $F_{b}-\Sigma_{b}=h^{p-1}\left(F_{b}^{(1)}-\Sigma_{b}^{(1)}\right)+\cdots+h\left(F_{b}^{(p-1)}-\Sigma_{b}^{(p-1)}\right).$
The characteristic polynomial of this matrix
(5.5) $\chi(F_{b}-\Sigma_{b})=\sum_{i=0}^{n}(-1)^{i}\mathrm{tr}\left(\Lambda^{i}(F_{b}-\Sigma_{b})\right)x^{n-i},$
where $\Lambda^{i}$ denotes the $i$th exterior product, satisfies the following properties: for each $1\leq i\leq n$,
@@
Since $F_{b}$ and $\Sigma_{b}$ both commute with the shift operators from 4.6 and 2.31, so is the difference $F_{b}-\Sigma_{b}$. Consider the shift operator $S_{\beta}$ for the cocharacter $\beta:S^{1}\to\mathbf{T}$ that vanishes on all $T$-weights and is weight $1$ for the $S^{1}$-action rotating the holomorphic symplectic form. By the invertibility of shift operators 3.24, we can conjugate $F_{b}-\Sigma_{b}$ by the shift operator which shows the invariance of the characteristic polynomial $\chi(F_{b}-\Sigma_{b})$ under shift of equivariant parameters $h\mapsto h-t$.
@@
For almost all $p$, the $p$-curvature and the quantum Steenrod operation for $b\in H^{2}_{\mathbf{T}}(X;\mathbb{F}_{p})$ agree.
@@
From 5.1, we know that the difference $F_{b}-\Sigma_{b}$ is a nilpotent operator acting on $H^{*}_{\mathbf{T}}(X;\Lambda)_{\mathrm{loc}}{[\![t,\theta]\!]}$ as a $H^{*}_{{\mathbb{Z}/p}\times\mathbf{T}}(\mathrm{pt})_{\mathrm{loc}}$-module. In particular, it suffices to show that $\Sigma_{b}$ and $F_{b}$ are simultaneously diagonalizable. From 2.20, $\Sigma_{b}$ commutes with $\nabla_{a}$ and hence with $F_{a}=\nabla_{a}^{p}-t^{p-1}\nabla_{a}$ for any $a\in H^{2}_{\mathbf{T}}(X;\mathbb{Z})$, and also $F_{b}$ and $F_{a}$ commute by the flatness of $\nabla$. In particular, it suffices to show that there exists $a\in H^{2}_{\mathbf{T}}(X;\mathbb{Z})$ for which $F_{a}$ has simple spectrum (i.e. diagonalizable with distinct eigenvalues) to show that $F_{a}=\Sigma_{a}$.
To check that the eigenvalues of $F_{a}$ are all distinct for some $a$, we consider the discriminant of $F_{a}$, which we denote by $D(F_{a})$. Since the equivariant parameter $t\in H^{2}_{\mathbb{Z}/p}(\mathrm{pt})$ is not inverted, one can consider the specialization $D(F_{a})|_{t=0}=D(a^{\ast_{\mathbf{T}}p}\ast_{\mathbf{T}})$; if the operator $a^{\ast_{\mathbf{T}}p}\ast_{\mathbf{T}}$ has simple spectrum, then the discriminant of it would be nonzero, and $D(F_{a})$ would also be nonzero. Therefore it suffices to check that $a^{\ast_{\mathbf{T}}p}\ast_{\mathbf{T}}$ has simple spectrum; this follows if $a{\ast_{\mathbf{T}}}$ has simple spectrum.
Choose $a\in\mathrm{Lie}(T)^{*}\cong H^{2}(X;\mathbb{Z})\subseteq H^{2}_{\mathbf{T}}(X;\mathbb{Z})$ which corresponds to a character $a:T\to S^{1}$. Consider the eigenvalues of $a\ast_{\mathbf{T}}|_{q^{A}=0}=a\cup_{\mathbf{T}}$; these are the weights of the (Borel–Weil) line bundle corresponding to $a$, restricted to $T$-fixed points of $X$. This line bundle is given by (the pullback under $T^{*}(G/B)\to G/B$ of the) associated line bundle $G\times^{B}\mathbb{C}\to G/B$, where we use the character $a:T\to S^{1}$ (and its unique extension to $B$) to make $\mathbb{C}$ into a $B$-representation.
At the point $B/B\in X^{T}$, this weight is tautologically given by $a$ itself, and at the other fixed points $wB/B\in X^{T}$, the weights are given by the Weyl group orbits $w\cdot a$ of $a$. When we choose $a$ generically (e.g. in the dominant Weyl chamber), these weights are indeed distinct in $H^{*}_{T}(\mathrm{pt};\mathbb{F}_{p})$ for $X=T^{*}(G/B)$ for all but finitely many primes $p$, as they lie in different Weyl chambers.
Therefore $a\cup_{\mathbf{T}}$ has simple spectrum, which implies that $a\ast_{\mathbf{T}}$ has simple spectrum. Hence $F_{a}$ has simple spectrum, and both $F_{b}$ and $\Sigma_{b}$ which commute with $F_{a}$ can be simultaneously diagonalized into the eigenbasis of $F_{a}$, so the desired result follows. ∎
@@
Proceeding as in Section 3.4, this result also applies to further examples including $X=T^{*}(G/P)$ (the generalized Springer resolutions), quiver varieties of type $A$, or hypertoric varieties, which have (i) isolated $X^{T}$ and hence a stable basis of homogeneous cohomological degree, (ii) (jointly) simple spectrum of the classical (hence also quantum) multiplication operator. The Springer resolution $T^{*}(G/B)$ in type $A$ are in the intersection of the first two classes of examples.